Find the length of side a in the right triangle, if b = 4 and c=5.
a =

Answer by richwmiller(17219)

3,4, 5 is a classic right triangle.
You should know it and 5,12,13
otherwise x^2+4^2=5^2
x^2+16=25
x^2=9
x=3
reject neg root